OpenStudy (anonymous):

how do you get the lcm of 10829 and 15015

OpenStudy (anonymous):

factor each number into a product of primes

OpenStudy (amistre64):

you dont ....

OpenStudy (anonymous):

\[10829=7^2\times 13\times 17\] \[15015=3 \times 5 \times 7 \times 11 \times 13\] so your lcm must be \[3\times 5\times 7^2\times 11\times 13\times 17\]

OpenStudy (amistre64):

isnt there a euler method for this?

OpenStudy (anonymous):

each prime that you see, so the highest power you see it in any one number

OpenStudy (anonymous):

this is the only method i know for finding the lcm besides starting at it and hoping it comes to you. that works fine for say 8 and 12, but not for these annoying numbers

OpenStudy (anonymous):

*staring

OpenStudy (amistre64):

lcm(a,b) = ab/gcd(a,b) gcd(a,b) = a/b = r b/r = s r/s = t repeat until x/y has no remainder

OpenStudy (amistre64):

15015/10829 = 1 R 46 10829/46 = 235 R 19 46/19 = 2 R 8 19/8 = 2 R 3 8/3 = 2 R 2 3/2 = 1 R 2 2/2 = 1 R 0 gcd(a,b) = 2 15015*10829/2 = .... well that would have been nice if i could remember it right :)

OpenStudy (anonymous):

looking for the greatest common divisor is what you are doing right? if i factored correctly it should be 7*13

OpenStudy (amistre64):

yeah